OpenStudy (anonymous):

The perimeter of a triangle is 36 cm and one of the legs is 3 cm larger than the other. Find the lengths of the three sides

OpenStudy (anonymous):

I am not sure we have enough information to do this. Has something been left out? Cleomenius.

OpenStudy (dumbcow):

a+b+c = 36 a = b+3 well yeah you need 1 more equation, it probably means this is a right triangle so use pythagorean theorem a^2 + b^2 = c^2

OpenStudy (anonymous):

Something is missing iandave..., I think som elittle info is left out!

OpenStudy (anonymous):

yes ofcourse dichiaraj u r right that dumb cow is right

OpenStudy (anonymous):

^ wrong thread ha ha

OpenStudy (anonymous):

hhahaha , dichiaraj well today i am first

OpenStudy (dumbcow):

so iandave, does it say if this is a right triangle? and do you need an answer or can you figure it out

OpenStudy (anonymous):

Insufficient data supplied. What type of triangle is it? Isosceles? Equilateral? What are some - or all - of the sizes of the angles?